q=500-0.5p and AC=q^2-59q+1315+200/q find the level of output and price at which profit will be maximized. find MC x=5 and interpret
1
Expert's answer
2020-05-29T09:27:32-0400
"Pr=TR-TC"


"TR=p\times q"


"TC=AC\times q"


"p=1000-2q"


"TR=1000q-2q^2"


"TC=q^3-59q^2+1315q+200"


"Pr=1000q-2q^2-q^3+59q^2-1315q-200"


"Pr=-q^3+57q^2-315q-200"


"Pr^\/=-3q^2+114q-315"

"-3q^2+114q-315=0"


"q^2-38q+105=0"


"q=35 or q=3"

q=3 is maximum point of function



"p=1000-2\times 3=994"


"MC=TC^\/"


"MC=3q^2-118q+1315"


"3q^2-118q+1315=3\times5^2-118\times5+1315=75-590+1315=800"

Under the given conditions, the maximum profit will not be.